I just wanted to know do I need to make corrections to my answers and are they right....?
7√100/√500 = 7√5/5
4/√21 = 4√21
√400/5 = 16√5/5
√10/81 = 10/9
-3√180h^4 = -3h√90
√144 = 12

Answer by stanbon(57347)

I need to make corrections to my answers and are they right....?
7√100/√500 =
= 7*10*10*sqrt(5)
= 700sqrt(5)
=======================
4/√21 = 4√21
OK
=======================
√(400/5)
= sqrt(80)
= sqrt(16)*sqrt(5)
= 4*sqrt(5)
------------------

√(10/81)
=sqrt(10)/9
------------------

-3√(180h^4)
= -3*sqrt(36h^2)*sqrt(5)
= -3*6h*sqrt(5)
= -18h*sqrt(5)
===================

√144 = 12
OK
=========================
Cheers,
Stan H.